Feed aggregator

APPROVED Replace meter in the 40 MVA

Recent Log Entries -

Log Number: 3855802Logbook(s): Body: 

All power fed from the 40MVA substation will be put onto the 33MVA substation Wednesday Morning beginning at ~7am.

Work has been coordinated by FML (powers/kujawa) with Operations (Suhring) and Cryo (Creel). FML has walked every loop, identified the break points, and have discussed the plan. FML is convinced there will be no interruption to power at any point in the process.

Once power and loads are removed from the 40MVA, the meter can be safely replaced.

Task Approved.

7-cell POFF optimized

Recent Log Entries -

Log Number: 3855788Logbook(s): Entry Makers: kiarraBody: 

7-cell was causing some issues this morning, it wouldn't stay in GDR mode, it kept
quenching. We kept walking down the gradient to see if that was the issue. Decided
to contact an expert -- Tomasz came in to take a look at things -- turned out the POFF
just needed to be optimized. It was off by ~10deg. 7-cell is back up and running!

MBHK101H and MHBM502H mismatch

Recent Log Entries -

Log Number: 3855784Logbook(s): Entry Makers: kiarraBody: 

MBHK101H current readback was very unstable -- tried reseating the card, no dice -- card was swapped.

MHBM502H was mismatched and wouldn't take setpoint -- tried reseating, no dice -- card was swapped.

Also having issues with MHBM904AH not responding -- it's possible the hardward in the cave isn't wired correctly. We may have to go into the cave to check things out. Will update when more information is available.

giun voltage drops to zero when magnet rack went down

Recent Log Entries -

Log Number: 3855781Logbook(s): Entry Makers: poelkerBody: 

Magnet rack ITF02 tripped off when Kiarra swapped a board, don't know why. But this took down the gun because the dipole bend magnet went to zero, or unresolved, or something. Ramping the gun back up. But at least this seemed to fix the first steering magnet, MBHK101H, now in agreement with setpoint request. We were able to turn the magnet rack back ON via epics

9.7 MeV/c beam to ITVM504

Recent Log Entries -

Log Number: 3855778Logbook(s): Entry Makers: poelkerBody: 

Tomasz coaxed the 7 cell back to life, hooray. With a small amount of steering associated with the inadvertant spot move, beam arrives at M504 viewer. Now time to lens center and quad center.

Don't know the exact beam momentum but GSets are the same as before. Beam bouncing on viewer M504, magnet mismatch issues, perhaps. Kiarra is swapping trim cards

EHS&Q REVIEW BSY and Halls Survey monument install & measure.

Recent Log Entries -

Log Number: 3855772Logbook(s): Body: 

Daily prior to working in the specific hall communicate with the hall work coordinator:
Hall A - Jessie Butler
Hall B - Denny Insley
Hall C - Walter Kellner
Hall D - Mark Stevens

Hall specific training & Physics Division Work Governance (SAF 116kd)
Hall A - SAF110
Hall B - SAF111
Hall C - SAF112
Hall D - SAF113

iocitfmag restored

Recent Log Entries -

Log Number: 3855771Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-10-12 07:34:26 The output from the myRestore execution was as follows: iocitfmag magnet_MAG_UITF@2020-10-12 07:34 188(188) restored Pause: 2 second(s) magnet_HVAL_UITF@2020-10-12 07:34 94(94) restored Pause: 2 second(s) magnet_LOOP_UITF@2020-10-12 07:34 94(94) restored Pause: 2 second(s) magnet_CMPQ_UITF@2020-10-12 07:34 188(188) restored system@2020-10-12 07:34 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ued -A -s -i iocitfmag (success) Not connected {MHDICESOLVOLTA

Pages

Subscribe to cebaf2013 aggregator